OpenStudy (anonymous):

solve for x: x-1^4=16

OpenStudy (anonymous):

\[x-1^{4}=16\]

OpenStudy (unklerhaukus):

\[1^4=1\times1\times1\times1\]

OpenStudy (anonymous):

i think x=2 but I'm not sure

OpenStudy (anonymous):

i mean x=3

OpenStudy (unklerhaukus):

\[1^4=1\times1\times1\times1=1\]

OpenStudy (unklerhaukus):

it this your question \[(x-1)^4=16\]?